if i have a problem like this
-3(x+5)=3(x-1) how would i slove it and check it the check part is what i dont understand how to do it

Found 3 solutions by doukungfoo, edjones, stanbon:
Answer by doukungfoo(195)

solving for x
-3(x+5)=3(x-1)
first multiply -3 across (x+5)
-3x - 15 = 3(x-1)
next multiply 3 across (x-1)
-3x - 15 = 3x - 3
now move all terms with x to the left side
-3x - 3x -15 = -3
Then move all terms without x to the right side
-3x - 3x = -3 + 15
combine like terms
-6x = 12
finally solve for x by dividing both sides by -6
-6x/-6 = 12/-6
x = -2
Ok so we have solved for x and found that x = -2
To check the answer just take the orginal equation
-3(x+5)=3(x-1)
and replace the x with -2
-3(-2+5)=3(-2-1)
now just do the math
-3(3) = 3(-3)
-9 = -9
we end up with a true state
so we know that -2 is the correct answer.

Answer by edjones(7311)

-3(x+5)=3(x-1)
-3x-15=3x-3
-6x=12 subtract 3x and add 15 to each side
x=-2 divide both sides by -6.
Check:
substitute -2 for x
-3(x+5)=3(x-1)
-3(-2+5)=3(-2-1)
-3*3=3*-3
-9=-9
Ed

Answer by stanbon(48568)

-3(x+5)=3(x-1)
-3x-15 = 3x-3
6x = -12
x = -2
---------------
Checking in the original equation:
-3(-2+5)= 3(-2-1)
-3(+3) = 3(-3)
-9 = -9
----------------
Cheers,
Stan H.